# Q-Mamba: On First Exploration of Vision Mamba for Image Quality Assessment

### v1

@@ 4.2.3 Analysis
LQMamba-S 8.66G 0.935 0.936 0.942 0.938 0.902 0.870 0.938 0.923 0.912 0.860 0.859 0.843 0.905
LQMamba-B 15.30G 0.941 0.939 0.944 0.941 0.916 0.889 0.939 0.927 0.910 0.859 0.870 0.857 0.911
Table 3: Performance comparison for transferable IQA.
  Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
    Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
    Without_tuning 0.535 0.501 0.726 0.744 0.789 0.784 0.726 0.744 0.553 0.549 0.595 0.550 0.600 0.612 0.675 0.649 0.636 0.587 0.667 0.637 0.460 0.425 0.815 0.838 0.642
Lin_Probe(R) 0.606 0.581 0.783 0.811 0.824 0.804 0.769 0.695 0.871 0.848 0.785 0.751 0.803 0.788 0.771 0.723 0.818 0.792 0.790 0.761 0.570 0.540 0.799 0.826 0.755
@@
0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
StylePrompt & R 0.921 0.913 0.944 0.945 0.877 0.853 0.906 0.847 0.931 0.912 0.889 0.860 0.876 0.857 0.905 0.849 0.929 0.909 0.899 0.859 0.918 0.911 0.926 0.927 0.898
4.3 Efficent transfer learning for Mamba-based IQA
In the context of transferable IQA tasks, we engaged in domain-specific training utilizing synthetic datasets (LIVE, KADID), authentic datasets (LIVEC, KonIQ), and AIGC datasets (AIGC2023, AGIQA3K). Each dataset underwent subsequent evaluations in the other two domains. We employed the StylePrompt technique, whereby we maintained the architecture’s integrity during cross-domain transfer by freezing it entirely, and solely fine-tuned the StylePrompt, utilizing merely $4\%$ of the total parameter volume. This approach achieved results that were nearly equivalent to those obtained through full-parameter training. The outcomes, as delineated in Table 3, present a comparative analysis of various fine-tuning strategies, conclusively validating the efficacy and efficiency of our proposed StylePrompt method in facilitating transferable IQA tasks.
4.4 Ablation Studies
4.4.1 Different scanning
Examining the outcomes from task-specific assessment in Table 1 and universal assessment in Table 2 , models equipped with local scanning have shown a definitive advantage over those utilizing cross-scanning techniques. This observation has led to an exploration of the effectiveness of local scanning methods.
4.4.2 Different model size
In our empirical analysis, as documented in Table 1 and Table 2, we discern that while QMamba-Base exhibits exceptionally robust quality perception capabilities, QMamba-Small either matches or exceeds the performance of QMamba-Base across the majority of the datasets. Although QMamba-Tiny displays a modest decline in performance metrics, it still delivers results that are competitive with current state-of-the-art (SOTA) methods, solely utilizing the capabilities of QMamba-Tiny.
Table 4: Ablation study on different prompt tuning strategies
@@ 4.4.2 Different model size
4.4.3 Different prompt tuning strategies
To validate the effectiveness of the StylePrompt Generation (SPG) process, we tested the direct learning of a set of affine parameters, $\gamma$ and $\beta$, to modulate the original features. The results in Table 4 demonstrate the outcomes of transfer learning using only the learned affine parameters, substantiating the effectiveness of our StylePrompt Generation phase.
Furthermore, to verify the efficacy of different prompt interaction methods within the StylePrompt, we evaluated various prompt interaction strategies with the original features. These included convolutional prompt interaction and cross-attention prompt interaction. Experimental results indicated that the interaction method employed by the StylePrompt with the original features achieves the best transfer results with minimal parameter usage. This further validates the effectiveness of our StylePrompt Injection (SPI) phase.
Table 5: Ablation study on the number of prompt
  Num of Prompt PLCC_Average SRCC_Average Average
  N=1 0.905 0.881 0.8928
N=2 0.905 0.882 0.8936
N=4 0.907 0.884 0.8951
N=6 0.911 0.890 0.9006
N=8 0.905 0.883 0.8939
N=10 0.906 0.883 0.8942
Table 6: Ablation study on the shape of prompts
  Prompt_shape(H,W) PLCC_Average SRCC_Average Average
  (1,1) 0.911 0.890 0.901
(7,7) 0.909 0.886 0.897
(14,14) 0.909 0.886 0.897
(28,28) 0.907 0.883 0.895
layer_HW 0.908 0.884 0.896
4.4.4 Different Prompt Designs
We investigated the impact of the number of prompt components on performance. Our findings revealed that setting the number of prompts to six yields optimal performance. Furthermore, we explored whether prompts with different spatial dimensions and sizes contribute to enhanced capabilities. Ultimately, we found that a spatial size of (1,1), which solely consists of the channel dimension, yields superior performance. We report the average results of these two ablation studies across multiple datasets in Tables 5 and Table 6. More detailed data will be provided in the appendix Table 8 and Table 9.
5 Conclusion
In this paper, we introduced Q-Mamba, a novel exploration of the Mamba model in the IQA field, which aims to enhance image quality assessment (IQA) from three perspectives: task-specific, universal, and transferable IQA. Our extensive experiments have demonstrated that Q-Mamba has great potential to outperform existing foundational models in IQA, including Swin Transformer, ViT, and CNNs, in both perceptual accuracy and computational efficiency. To improve the perception transferability of Q-Mamba, we developed StylePrompt, a powerful tuning mechanism, to enable the effective perception knowledge transfer of Q-Mamba to various downstream IQA tasks even with large domain shifts, by exploiting only extra $4\%$ of whole parameters to adjust the statistics of perception features of Q-Mamba. We believe these findings can prompt the further development of IQA and pave the way for future research of Mamba-based quality assessment.
Appendix A Appendix / supplemental material
A.1 Additional experimental results
A.1.1 Task-specific results on AIGCIQA
Table 7: Results of AIGC on various backbones
  Model AGIQA3K AIGCIQA2023
  FLOPs PLCC SRCC PLCC SRCC PLCC_Average SRCC_Average Average
     ResNet-50(23.51M) 4.11G 0.901 0.840 0.795 0.797 0.848 0.819 0.833
ResNet-101(42.50M) 7.83G 0.907 0.847 0.834 0.831 0.871 0.839 0.855
ResNet-152(58.15M) 11.53G 0.901 0.832 0.841 0.834 0.871 0.833 0.852
ViT-T(5.52M) 1.26G 0.865 0.787 0.760 0.766 0.813 0.777 0.795
ViT-S(21.67M) 4.61G 0.891 0.819 0.842 0.822 0.867 0.821 0.844
ViT-B(85.80M) 17.58G 0.897 0.830 0.853 0.835 0.875 0.833 0.854
Swin-T(27.52M) 4.51G 0.906 0.847 0.867 0.844 0.887 0.846 0.866
Swin-S(48.84M) 8.77G 0.908 0.849 0.875 0.857 0.892 0.853 0.872
Swin-B(86.74M) 15.47G 0.909 0.852 0.886 0.863 0.898 0.858 0.878
QMamba-T (27.99M) 4.47G 0.913 0.858 0.888 0.873 0.901 0.866 0.883
QMamba-S (49.37M) 8.71G 0.912 0.858 0.889 0.875 0.901 0.867 0.884
QMamba-B (87.53M) 15.35G 0.914 0.861 0.886 0.868 0.900 0.865 0.882
LQMamba-T(29.87M) 4.44G 0.914 0.862 0.884 0.868 0.899 0.865 0.882
LQMamba-S(52.91M) 8.66G 0.913 0.864 0.888 0.869 0.901 0.867 0.884
LQMamba-B(93.79M) 15.30G 0.915 0.858 0.888 0.871 0.902 0.865 0.883
A.1.2 Detailed Results of Ablation Studies
We report the detailed data of our experiments on Prompt design in Table 8 and Table 9.
Table 8: Ablation study results for different numbers of prompts (N).
@@ A.1.2 Detailed Results of Ablation Studies
Table 9: Ablation study for different prompt shape
@@
A.2 Limitations
Q-Mamba heavily relies on pre-trained weights from ImageNet-1K, which may limit its applicability to domains with significantly different data distributions. Future work could explore pre-training on more diverse datasets to improve generalization capabilities. Despite the efficiency improvements brought by the StylePrompt, the computational demands for training and deploying Q-Mamba on very large-scale datasets or in real-time applications might still be substantial. Investigating methods to further reduce computational complexity without sacrificing performance could be beneficial. The current study focuses on specific types of distortions, and there may be other types of distortions that have not been sufficiently explored. Expanding the evaluation to cover a broader range of distortions could provide a more comprehensive validation of Q-Mamba’s robustness. Given that synthetic datasets might not fully capture the complexity and variability of authentic data, there is a risk of overfitting to these synthetic examples. Further evaluations on more diverse and extensive authentic datasets would help ensure the model’s robustness and practical applicability. While Q-Mamba shows promising results in cross-domain transferability, its effectiveness across vastly different domains (e.g., medical imaging versus natural images) has not been thoroughly validated. Further studies are needed to test and possibly adapt Q-Mamba for such diverse applications.
A.3 Broader Impacts
Improving image quality assessment (IQA) through Q-Mamba allows users across various platforms to enjoy higher quality images, enhancing user experience in social media, streaming services, and other digital platforms, thus leading to greater satisfaction and engagement. The enhancement of image quality can significantly benefit individuals with visual impairments, as better image clarity and quality make content more accessible, thereby improving the inclusivity of digital media. Additionally, the computational efficiency of Q-Mamba can result in more efficient image processing workflows, reducing the time and energy required for image enhancement, compression, and transmission. This can lead to cost savings and environmental benefits due to lower energy consumption. Moreover, the transferability of Q-Mamba across different domains can improve image quality in fields such as medical imaging, satellite imagery, and autonomous driving, thereby enhancing the accuracy and reliability of critical applications.
While Q-Mamba presents these positive societal impacts, it is important to acknowledge and mitigate potential negative effects. Enhancing image quality and clarity may inadvertently lead to privacy concerns, as higher quality images can capture more detailed information about individuals and environments. However, our work adheres to ethical guidelines and uses publicly available, open-source datasets, ensuring compliance with various constraints and regulations. By focusing on responsible and ethical use, we aim to maximize the benefits of Q-Mamba while minimizing any adverse societal impacts.

### v2

@@ A.2.2 Universal IQA detailed data
Table 6: Performance comparison for universal IQA.
A.2.3 Transferable IQA results trained on the AIGC domain
We provide the detailed results of Transferable IQA trained in the AIGC domain in Table 7.
Train AIGC2023 & AGIQA3K (AIGC)
Test KonIQ LIVEC KADID LIVE
Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
DEIQT 0.578 0.505 0.684 0.638 0.498 0.464 0.752 0.750 0.601
LoDa 0.564 0.500 0.542 0.548 0.555 0.524 0.840 0.833 0.622
Without_tuning 0.636 0.587 0.667 0.637 0.460 0.425 0.815 0.838 0.642
Lin_Probe(R) 0.818 0.792 0.790 0.761 0.570 0.540 0.799 0.826 0.755
@@ A.2.3 Transferable IQA results trained on the AIGC domain
0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
StylePrompt & R 0.929 0.909 0.899 0.859 0.918 0.911 0.926 0.927 0.898
Table 7: Performance comparison for transferable IQA (trained on AIGC).
A.2.4 Detailed results of ablation studies
  • Ablation study on different prompt tuning strategies.
    We have provided the detailed experimental data for the ablation study on different prompt tuning strategies in Table 8.
  • Ablation study results for different prompt designs.
    We investigated the impact of the number of prompt components on performance and found that setting the number to six yields optimal results. Additionally, we explored whether varying the spatial dimensions and sizes of prompts would enhance performance. Our findings show that a spatial size of (1,1), focusing solely on the channel dimension, offers the best results. The outcomes of these ablation studies across multiple datasets are presented in Tables 9 and 10.
@@ A.2.4 Detailed results of ablation studies
Table 8: Ablation study on different prompt tuning strategies
@@
Table 9: Ablation study results for different numbers of prompts (N).
@@
Table 10: Ablation study for different prompt shape
A.3 Limitations
Q-Mamba heavily relies on pre-trained weights from ImageNet-1K, which may limit its applicability to domains with significantly different data distributions. Future work could explore pre-training on more diverse datasets to improve generalization capabilities. Despite the efficiency improvements brought by the StylePrompt, the computational demands for training and deploying QMamba on very large-scale datasets or in real-time applications might still be substantial. Investigating methods to further reduce computational complexity without sacrificing performance could be beneficial. The current study focuses on specific types of distortions, and there may be other types of distortions that have not been sufficiently explored. Expanding the evaluation to cover a broader range of distortions could provide a more comprehensive validation of QMamba’s robustness. Given that synthetic datasets might not fully capture the complexity and variability of authentic data, there is a risk of overfitting to these synthetic examples. Further evaluations on more diverse and extensive authentic datasets would help ensure the model’s robustness and practical applicability. While QMamba shows promising results in cross-domain transferability, its effectiveness across vastly different domains (e.g., medical imaging versus natural images) has not been thoroughly validated. Further studies are needed to test and possibly adapt QMamba for such diverse applications.